# The super learner for time-to-event outcomes: A tutorial

### v1

@@ 3.3 SL step 4: Estimating the mean expected loss
As noted above, a loss function is a measure of the difference between estimates of the target estimand and the true value. The loss function is defined at a individual level. When the focus is on estimating the conditional survival probability it is natural to consider the squared difference between $S(\tau|X=x)$ and the event indicator $I(T>\tau)$ as the loss function: i.e. $\left(I(T>\tau)-S(\tau|X=x)\right)^{2}$. However, when there is right censoring we do not observe $I(T>\tau)$ for everyone in the data, and instead we observe $I(\tilde{T}>\tau)$ and $\Delta$. To account for censoring, a loss function that uses an inverse probability of censoring weight (IPCW) is defined as
@@
This is equivalent to individual contributions to the Brier score. Using this loss function requires an estimate of the conditional survival probability for censoring $G(\cdot|X)$, and we discuss this further below. Polley and van der Laan (2011) also considered alternative loss functions, which have the major advantage of avoiding having to estimate the conditional censoring distribution, which is a nuisance parameter in (1). The alternative loss functions are based on the hazard rather than the survival probability. One is a squared error loss summarised across all times:
@@
Polley and van der Laan (2011) comment that all three of the above loss functions are valid in this case, but that (1) is the most ‘natural’ to use as it directly targets the target estimand, which is the conditional survival probability.
After fitting the candidate learners, the next step is to estimate the expected loss in each of the $K$ data folds based on the chosen loss function. To estimate the expected loss based on $L_{IPCW}$ in (1) requires an estimate of $G(\tilde{T}|X)$ or $G(\tau|X)$ (for individuals who remain under observation at time $\tau$). If the censoring is assumed to be independent then $G(t|X)$ can be replaced by the marginal probability $G(t)$, which can be estimated non-parametrically using the Kaplan-Meier estimator, and this is the approach taken by Polley and van der Laan (2011). The censoring weights could alternatively be allowed to depend on $X$, for example through use of SL to estimate $G(t|X)$. This would be performed separately from the main SL. The censoring probabilities are estimated in the validation data, i.e. using $D_{k}$. Assuming independent censoring, the mean expected loss for candidate learner $j$ in validation set $D_{k}$ is then estimated using
@@
and the expected loss based on (3) is estimated similarly. The estimated mean expected loss for each candidate learner $j$ is then given by $\bar{L}^{(j)}=\sum_{k=1}^{K}\hat{L}^{(j,k)}$, $j=1,\ldots,p$.
@@ 3.4 SL step 5: Deriving an ensemble
The non-ensemble SL chooses the candidate learner with the best (i.e. lowest) cross-validated mean expected loss based on the chosen loss function. In this section we focus on the ensemble SL, which derives an optimally weighted combination of the candidate learners, in such a way that the combination minimizes the chosen loss function. To give the general idea, first consider the simplified setting of a continuous outcome $Y$ and suppose we had fitted $p$ candidate learners resulting in cross-validated fitted values (conditional expectations) $\hat{Y}_{1},\ldots,\hat{Y}_{p}$ for each individual in the data. The SL ensemble would then be derived based on the model $E(Y|\hat{Y}_{1},\ldots,\hat{Y}_{p})=\alpha_{1}\hat{Y}_{1}+\cdots+\alpha_{p}\hat{Y}_{p}$, with the parameters $\alpha_{j}$ ($j=1,\ldots,p$) estimated subject to the constraints that $\alpha_{j}\geq 0$ for all $j=1,\ldots,p$, and $\sum_{j=1}^{p}\alpha_{j}=1$. This can be achieved by fitting the regression model using non-negative least squares and then scaling the coefficients to sum to 1. This minimises the mean squared error, which is the appropriate loss function for a continuous outcome.
For the time-to-event setting different choices for the loss function were summarised above, and these correspond to slightly different regressions to derive the SL ensemble. We begin by considering a SL based on the loss function $L_{L2}$, as this is most similar to the simpler case of a continuous outcome outlined above. The aim is to obtain an algorithm from which we can obtain estimates of the conditional hazard $Q(t|X)$ in such a way that $L_{L2}$ is minimized. This can be achieved by a regression of $\Delta_{i}(t)$ on the estimates of the conditional hazard from the $p$ candidate learners, $\psi_{i,j}(t)=\hat{Q}_{j}(t|X=x_{i})$ for $j=1,\ldots,p$. The model is fitted across all times $t=0,\ldots,\tau-1$ combined using the long data set. The regression equation is
$E(\Delta(t)|\psi_{1}(t),\ldots,\psi_{p}(t))=\sum_{j=1}^{p}\alpha_{j}\psi_{j}(t)$ (6)
@@
If the chosen loss function is $L_{loglik}$ then a modified version of the above procedure is needed to derive a SL such that $L_{loglik}$ is minimized. Polley and van der Laan (2011) recommend transforming the conditional hazards using $g(\psi_{i,j}(t))=\mathrm{logit}\left\{\psi_{i,j}(t)\right\}$, and using the logistic regression equation
@@
They note that a constraint on the coefficients $\alpha_{j}$ ($j=1,\ldots,p$) is not required under this approach because the transformed conditional hazards $g(\psi_{i,j})$ are not constrained. The SL estimator for the conditional hazard based on $L_{loglik}$ is then
@@
Lastly we consider the case where the chosen loss function is $L_{IPCW}$, which is based on the conditional survival probability rather than the conditional hazard. Let $\phi_{i,j}=\hat{S}_{j}(\tau|X=x_{i})$ for $j=1,\ldots,p$. The SL to minimize this loss function is derived based on the regression equation
$E(I(T>\tau)|\phi_{1},\ldots,\phi_{p})=\sum_{j=1}^{p}\alpha_{j}\phi_{i,j}.$ (10)
@@
The regression in equation (10) is fitted using a data set with one row per individual, because $\phi_{i,j}$ refers to the survival probability at the specified time $\tau$ only. This differs from the fitting of the regressions based on $L_{L2}$ and $L_{loglik}$, which are fitted on the long data.
@@ 4.1 Overview
The discrete-time SL has some disadvantages. The discretisation of follow-up time means that we do not make full use of the information on event times that are in the data. Flexible modelling of time as a covariate in candidate learners for the baseline hazard is also a challenge. The two-step approach recommended by Polley and van der Laan (2011) is not implemented in the main R package SuperLearner available for implementation. The requirement to discretise the data, resulting in a data set with many rows per individual, can also bring challenges in terms of computing time and power.
In this section we describe two continuous-time SL approaches. Westling and others (2023) described a SL for continuous-time data (or a mixture of continuous and discrete data). They are motivated by the need to estimate conditional survival and censoring distributions, $S(t|X)$ and $G(t|X)$, because they feature as nuisance functions in a wider doubly-robust estimation procedure aiming to estimate survival curves under different treatments. We focus on their SL for estimation of target estimand $S(t|X)$. Munch and Gerds (2024) described a SL for continuous event time data, referred to as the ‘state learner’. They focus on use of a SL to estimate state occupation probabilities in a setting that can include more than one event type of interest, for example where there are competing events. We focus on the use of their approach when there is just a single absorbing event type of interest (e.g. death). The method of Westling and others (2023) includes an ensemble step, whereas the ‘state learner’ of Munch and Gerds (2024) is a non-ensemble SL. Both methods involve loss functions that require estimates of the conditional censoring distribution $G(t|X)$, but they enable this to be done in a flexible way, making use of a SL for the conditional censoring distribution. The estimation of the conditional censoring distribution is integrated within the same SL procedure as conditional survival distribution.
@@ 4.2.1 SL steps 1-3: Cross-validation
This method targets estimation of the whole conditional survival distribution $S(t|X)$, for $0<t\leq\tau$ say. Any model that gives rise to an estimate of $S(t|X)$ for any value of $t$ (within a specified range) could be a candidate for inclusion in the SL. In their example application Westling and others (2023) include the following candidate learners: parametric survival models, semiparametric proportional hazard models, generalized additive Cox models, and random survival forests. Each candidate learner $j=1,\ldots,p$ is fitted in training set $D_{-k}$ and is then used to obtain estimates $\hat{S}_{j}(t|X=x_{i})$ for individuals in validation set $D_{k}$. In order to evaluate predictive performance of candidate learners (see step 4), this approach requires estimates $\hat{S}_{j}(t|X=x_{i})$ on a grid of times. We emphasise that this is different from the discretisation of time for estimation of $S(\tau|X)$ in the discrete-time SL. Suppose that we divide the time interval of interest $(0,\tau)$ into a grid $V$ of evenly spaced times with differences $v$ (i.e. $V=\{v,2v,3v,\ldots,\tau-v,\tau\}$). Cross-validated estimates of $\hat{S}_{j}(t|X=x_{i})$ are obtained at each time $t$ in $V$. Using each candidate learner $j$ we can then create a vector of estimated conditional survival probabilities for each person $i$ in $D$ at each time point on the grid $V$: $\hat{S}_{i,j}^{V}=\left(\hat{S}_{j}(v|X=x_{i}),\hat{S}_{j}(2v|X=x_{i}),\ldots,\hat{S}_{j}(\tau|X=x_{i})\right)$, for $i=1,\ldots,n$ and $j=1,\ldots,p$.
@@ 4.2.2 SL Step 4: Estimating the mean expected loss
where $f_{G}(t)=1-\Delta I(\tilde{T}\leq t)/G(\tilde{T}|X)$. This loss function is a weighted integrated Brier score. Their approach can be used under a pre-specified model for the censoring distribution $G(t|X)$, or $G(t|X)$ can itself be estimated using a SL using an approach that combines SLs for both $S(t|X)$ and $G(t|X)$. We first outline the use of the approach of Westling and others (2023) under a pre-specified model for $G(t|X)$, before outlining their extension to using SLs for both $S(t|X)$ and $G(t|X)$ (Section 4.2.4). Suppose that a model for the conditional survival distribution $G(t|X)$ has been pre-specified, for example as a Cox model or random survival forest, and let $\hat{G}(\tilde{T}|X)$ denote the resulting cross-validated estimates. Based on $L_{S,G}$, the expected loss for candidate learner $j$ in fold $k$ is obtained as follows, where the integral is approximated via a sum using the grid of times $V$:
@@
As before, we then obtain the estimated mean expected loss $\bar{L}_{j}$ for each candidate learner $j=1,\ldots,p$ using the average of $\hat{L}_{S,G}^{(j,k)}$ over the $K$ folds (step 4 in Section 2.2). As before, the best performing individual learner, i.e. the one with smallest cross-validated risk is the non-ensemble SL.
@@ 4.2.3 SL step 5: Deriving an ensemble
Derivation of the SL ensemble makes use of the cross-validated estimated survival probabilities $\hat{S}_{i,j}^{V}$ ($j=1,\ldots,p$), and the aim is to derive a weighted combination of the estimates from the $p$ candidate learners such that the loss function $L_{S,G}$ in (12) is minimised. The loss function features $f_{G}(t)=1-\Delta I(\tilde{T}\leq t)/G(\tilde{T}|X)$ and we obtain $f_{G}(i,t)=\left(1-\frac{\Delta_{i}I(\tilde{T}_{i}\leq t)}{\hat{G}(\tilde{T}_{i}|X=x_{i})}\right)$ for each person at each time point $t$ in the grid $V$, giving the outcome vector $f^{V}_{G}(i)=\left(f_{G}(i,v),f_{G}(i,2v),\ldots,f_{G}(i,\tau)\right)$.
The ensemble SL is then derived by performing a non-negative least squares linear regression of $f^{V}_{G}(i)$ on $\hat{S}_{i,1}^{V},\ldots,\hat{S}_{i,p}^{V}$ across all times in $V$ combined, i.e. a model with linear predictor $\sum_{j=1}^{p}\alpha_{j}\hat{S}_{j}(v|X=x_{i})$, where the $\alpha_{j}$ are constrained to be non-negative. After fitting the model the estimates $\alpha_{1},\ldots,\alpha_{p}$ are scaled to sum to 1 as described earlier. The resulting SL estimator is of the form
@@
4.2.4 Extending to estimate the censoring distribution via SL
We now outline how the method extends to allow for using a SL for the censoring distribution $G(t|X)$ as well as for $S(t|X)$. This requires candidate learners and loss functions for both distributions. Westling and others (2023) specify the following loss function for the conditional censoring distribution over the interval $(0,\tau)$, which is analogous to the loss function $L_{S,G}$ for the survival distribution in (12):

### v2

@@ 3.3 SL step 4: Estimating the mean expected loss
After fitting the candidate learners, the next step is to estimate their predictive performance based on a chosen loss function. As noted in Section 2.2, a loss function is a measure of the difference between estimates of the target estimand and the true value. The loss function is defined at a individual level. When the focus is on estimating the conditional survival probability it is natural to consider the squared difference between $S(\tau|X=x)$ and the event indicator $I(T>\tau)$ as the loss function: i.e. $\left(I(T>\tau)-S(\tau|X=x)\right)^{2}$. However, when there is right censoring we do not observe $I(T>\tau)$ for everyone in the data, and instead we observe $I(\tilde{T}>\tau)$ and $\Delta$. To account for censoring, a loss function that uses an inverse probability of censoring weight (IPCW) is defined as
@@
This is equivalent to individual contributions to the Brier score (Graf and others, 1999). Using this loss function requires an estimate of the conditional survival probability for censoring $G(\cdot|X)$, and we discuss this further below. Polley and van der Laan (2011) also considered alternative loss functions, which have the advantage of avoiding having to estimate the conditional censoring distribution, which is a nuisance parameter in (1). The alternative loss functions are based on the hazard rather than the survival probability. One is a squared error loss summarised across all times:
@@
Polley and van der Laan (2011) comment that all three of the above loss functions are valid, but that (1) is the most ‘natural’ to use as it directly targets the target estimand, which is the conditional survival probability.
For each candidate learner, we estimate the expected loss in each of the $K$ data folds based on the chosen loss function. To estimate the expected loss based on $L_{IPCW}$ in (1) requires an estimate of $G(\tilde{T}|X)$ or $G(\tau|X)$ (for individuals who remain under observation at time $\tau$). If the censoring is assumed to be independent then $G(t|X)$ can be replaced by the marginal probability $G(t)$, which can be estimated non-parametrically using the Kaplan-Meier estimator, and this is the approach taken by Polley and van der Laan (2011). The censoring weights could alternatively be allowed to depend on $X$, for example through use of a SL to estimate $G(t|X)$. This would be performed separately from the main SL. The censoring probabilities are estimated in the validation data, i.e. using $D_{k}$. Assuming independent censoring, the expected loss for candidate learner $j$ in validation set $D_{k}$ is then estimated using
@@
and the expected loss based on (3) is estimated similarly. The estimated mean expected loss for each candidate learner $j$ is then obtained by taking the mean across the $K$ folds: $\bar{L}^{(j)}=\frac{1}{K}\sum_{k=1}^{K}\hat{L}^{(j,k)}$, $j=1,\ldots,p$.
@@ 3.4 SL step 5: Deriving an ensemble
The non-ensemble SL chooses the candidate learner with the best (i.e. lowest) cross-validated mean expected loss, based on the chosen loss function. In this section we focus on the ensemble SL, which derives an optimally weighted combination of the candidate learners, in such a way that the combination minimizes the chosen loss function. To give the general idea, first consider the simplified setting of a continuous outcome $Y$ and suppose we had fitted $p$ candidate learners resulting in cross-validated fitted values (conditional expectations) $\hat{Y}_{1},\ldots,\hat{Y}_{p}$ for each individual in the data. The SL ensemble would then be derived based on the model $E(Y|\hat{Y}_{1},\ldots,\hat{Y}_{p})=\alpha_{1}\hat{Y}_{1}+\cdots+\alpha_{p}\hat{Y}_{p}$, with the parameters $\alpha_{j}$ ($j=1,\ldots,p$) estimated subject to the constraints that $\alpha_{j}\geq 0$ for all $j=1,\ldots,p$, and $\sum_{j=1}^{p}\alpha_{j}=1$. This can be achieved by fitting the regression model using non-negative least squares and then scaling the coefficients to sum to 1. This minimises the mean squared error, which is the appropriate loss function for a continuous outcome.
For the time-to-event setting different choices for the loss function were summarised above, and these correspond to slightly different regressions to derive the SL ensemble. We begin by considering a SL based on the loss function $L_{L2}$ in (2), as this is most similar to the simpler case of a continuous outcome outlined above. The aim is to obtain an algorithm from which we can obtain estimates of the conditional hazard $Q(t|X)$ in such a way that $L_{L2}$ is minimized. This can be achieved by a regression of $\Delta_{i}(t)$ on the estimates of the conditional hazard from the $p$ candidate learners, $\psi_{i,j}(t)=\hat{Q}_{j}(t|X=x_{i})$ for $j=1,\ldots,p$. The model is fitted across all times $t=0,\ldots,\tau-1$ combined using the long data set. The regression equation is
$E\{\Delta(t)|\psi_{1}(t),\ldots,\psi_{p}(t)\}=\sum_{j=1}^{p}\alpha_{j}\psi_{j}(t)$ (6)
@@
If the chosen loss function is $L_{loglik}$ (equation (3)) then a modified version of the above procedure is needed to derive a SL such that $L_{loglik}$ is minimized. Polley and van der Laan (2011) recommend transforming the conditional hazards using $g(\psi_{j}(t))=\mathrm{logit}\left\{\psi_{j}(t)\right\}$, and using the logistic regression equation
@@
They note that a constraint on the coefficients $\alpha_{j}$ ($j=1,\ldots,p$) is not required under this approach because the transformed conditional hazards $g(\psi_{j})$ are not constrained. The SL estimator for the conditional hazard based on $L_{loglik}$ is then
@@
Lastly we consider the case where the chosen loss function is $L_{IPCW}$ (equation (1)), which is based on the conditional survival probability rather than the conditional hazard. Let $\phi_{j}=\hat{S}_{j}(\tau|X=x)$ for $j=1,\ldots,p$. The SL to minimize this loss function is derived based on the regression equation
$E\{I(T>\tau)|\phi_{1},\ldots,\phi_{p}\}=\sum_{j=1}^{p}\alpha_{j}\phi_{j}.$ (10)
@@
The regression in equation (10) is fitted using a data set with one row per individual, because $\phi_{j}$ refers to the survival probability at the specified time $\tau$ only. This differs from the fitting of the regressions based on $L_{L2}$ and $L_{loglik}$, which are fitted on the long data.
@@ 4.1 Overview
The discrete-time SL has some disadvantages. The discretisation of follow-up time means that we do not make full use of the information on event times that are in the data. Flexible modelling of time as a covariate in candidate learners for the baseline hazard is also a challenge. The two-step approach recommended by Polley and van der Laan (2011) is not implemented in the main R package SuperLearner available for implementation (see Section 5). The requirement to discretise the data, resulting in a data set with many rows per individual, can also bring computational challenges.
In this section we describe two continuous-time SL approaches, i.e. approaches that do not require discretisation of follow-up time. Westling and others (2023) described a SL for continuous-time data (or a mixture of continuous- and discrete-time data). They are motivated by the need to estimate conditional survival and censoring distributions, $S(t|X)$ and $G(t|X)$, because they feature as nuisance functions in a wider causal estimation procedure aiming to estimate survival curves under different treatments using a doubly-robust approach. We focus on their SL for estimation of target estimand $S(t|X)$. Another continuous-time SL approach was proposed by Munch and Gerds (2025), who referred to their approach as the ‘joint survival super learner’ (in an earlier version of their paper it was referred to as the ‘state learner’ Munch and Gerds (2024)). They focus on use of a SL to estimate state occupation probabilities in a setting that can include more than one event type of interest, for example where there are competing events. We focus on the use of their approach when there is just a single absorbing event type of interest (e.g. death). The method of Westling and others (2023) includes an ensemble step, whereas the joint survival SL of Munch and Gerds (2025) is a non-ensemble SL. Both methods involve loss functions that require estimates of the conditional censoring distribution $G(t|X)$, but they enable this to be done in a flexible way, making use of a SL for the conditional censoring distribution.
@@ 4.2.1 SL steps 1-3: Cross-validation
This method targets estimation of the whole conditional survival distribution $S(t|X)$, for $0<t\leq\tau$ say. Any model that gives rise to an estimate of $S(t|X)$ for any value of $t$ (within a specified range) could be a candidate learner for inclusion in the SL. In their example application Westling and others (2023) include the following candidate learners: parametric survival models, semiparametric proportional hazard models, generalized additive Cox models, and random survival forests. Each candidate learner $j=1,\ldots,p$ is fitted in training set $D_{-k}$ and is then used to obtain estimates $\hat{S}_{j}(t|X=x_{i})$ for individuals in validation set $D_{k}$. In order to evaluate predictive performance of candidate learners (see step 4), this approach requires estimates $\hat{S}_{j}(t|X=x_{i})$ for each $t$ on a grid of times. Suppose that we divide the time interval of interest $(0,\tau)$ into a grid $V$ of evenly spaced times with differences $v$ (i.e. $V=\{v,2v,3v,\ldots,\tau-v,\tau\}$). Cross-validated estimates of $\hat{S}_{j}(t|X=x_{i})$ are obtained at each time $t$ in $V$, i.e. $\hat{S}_{j}(v|X=x_{i}),\hat{S}_{j}(2v|X=x_{i}),\ldots,\hat{S}_{j}(\tau|X=x_{i})$. Using each candidate learner $j$ we can then create a vector of estimated conditional survival probabilities for each person $i$ in $D$ at each time point on the grid $V$: $\hat{S}_{i,j}^{V}=\left(\hat{S}_{j}(v|X=x_{i}),\hat{S}_{j}(2v|X=x_{i}),\ldots,\hat{S}_{j}(\tau|X=x_{i})\right)$, for $i=1,\ldots,n$ and $j=1,\ldots,p$. We emphasise that the use of the time grid here is different from the discretisation of time for estimation of $S(\tau|X)$ in the discrete-time SL. Using a fine time grid is recommended here, as the grid is used in the next step to estimate an integral.
@@ 4.2.2 SL Step 4: Estimating the mean expected loss
where $f_{G}(t)=1-\Delta I(\tilde{T}\leq t)/G(\tilde{T}|X)$. This loss function is a weighted integrated Brier score. It can be used under a pre-specified model for the censoring distribution $G(t|X)$, or $G(t|X)$ can itself be estimated using a SL using an approach that combines SLs for both $S(t|X)$ and $G(t|X)$. We first outline the use of the approach of Westling and others (2023) under a pre-specified model for $G(t|X)$, before outlining their extension to using SLs for both $S(t|X)$ and $G(t|X)$ (Section 4.2.4). Suppose that a model for the conditional survival distribution $G(t|X)$ has been pre-specified, for example as a Cox model or random survival forest, and let $\hat{G}(\tilde{T}|X)$ denote the resulting cross-validated estimates. Based on $L_{S,G}$, the expected loss for candidate learner $j$ in fold $k$ is obtained as follows, where the integral in (12) is approximated via a sum using the grid of times $V$:
@@
As before, we then obtain the estimated mean expected loss $\bar{L}^{(j)}$ for candidate learner $j$ ($j=1,\ldots,p$) using the average of $\hat{L}_{S,G}^{(j,k)}$ over the $K$ folds (step 4 in Section 2.2). As before, the best performing individual learner (the one with smallest mean expected loss) is the non-ensemble SL.
@@ 4.2.3 SL step 5: Deriving an ensemble
Derivation of the SL ensemble makes use of the cross-validated estimated survival probabilities on the time grid $V$ for the $p$ candidate learners, $\hat{S}_{i,j}^{V}$ ($j=1,\ldots,p$). The aim is to derive a weighted combination of the estimates from the $p$ candidate learners such that the loss function $L_{S,G}$ in (12) is minimised. The loss function involves $f_{G}(t)=1-\Delta I(\tilde{T}\leq t)/G(\tilde{T}|X)$ and we can obtain the estimates $\hat{f}_{G}(i,t)=\left(1-\frac{\Delta_{i}I(\tilde{T}_{i}\leq t)}{\hat{G}(\tilde{T}_{i}|X=x_{i})}\right)$ for each person at each time point $t$ in the grid $V$, giving the outcome vector $\hat{f}^{V}_{G}(i)=\left(\hat{f}_{G}(i,v),\hat{f}_{G}(i,2v),\ldots,\hat{f}_{G}(i,\tau)\right)$. The ensemble SL is then derived by performing a non-negative least squares linear regression of $\hat{f}^{V}_{G}(i)$ on $\hat{S}_{i,1}^{V},\ldots,\hat{S}_{i,p}^{V}$ across all times in $V$ combined, i.e. a model with linear predictor $\sum_{j=1}^{p}\alpha_{j}\hat{S}_{j}(t|X=x_{i})$, where the $\alpha_{j}$ are constrained to be non-negative. The model is fitted on a ‘long’ data set which includes the estimates $\hat{f}_{G}(i,t)$ and $\hat{S}_{j}(t|X=x_{i})$ for each person $i$ in $D$ at each time $t$ on the grid $V$. After fitting the model the estimates of $\alpha_{1},\ldots,\alpha_{p}$ are scaled to sum to 1 as described earlier. The resulting SL estimator is of the form
@@
4.2.4 Extending to estimate the censoring distribution via super learner
We now outline how this method extends to allow for using a SL for the censoring distribution $G(t|X)$ as well as for $S(t|X)$. This requires candidate learners and loss functions for both distributions. Westling and others (2023) specify the following loss function for the conditional censoring distribution over the interval $(0,\tau)$, which is analogous to the loss function $L_{S,G}$ for the survival distribution in (12):